# Integrating Intermediate Layer Optimization and Projected Gradient Descent for Solving Inverse Problems with Diffusion Models

### v2

@@ Abstract
Inverse problems (IPs) involve reconstructing signals from noisy observations. Recently, diffusion models (DMs) have emerged as a powerful framework for solving IPs, achieving remarkable reconstruction performance. However, existing DM-based methods frequently encounter issues such as heavy computational demands and suboptimal convergence. In this work, building upon the idea of the recent work DMPlug (Wang et al., 2024), we propose two novel methods, DMILO and DMILO-PGD, to address these challenges. Our first method, DMILO, employs intermediate layer optimization (ILO) to alleviate the memory burden inherent in DMPlug. Additionally, by introducing sparse deviations, we expand the range of DMs, enabling the exploration of underlying signals that may lie outside the range of the diffusion model. We further propose DMILO-PGD, which integrates ILO with projected gradient descent (PGD), thereby reducing the risk of suboptimal convergence. We provide an intuitive theoretical analysis of our approaches under appropriate conditions and validate their superiority through extensive experiments on diverse image datasets, encompassing both linear and nonlinear IPs. Our results demonstrate significant performance gains over state-of-the-art methods, highlighting the effectiveness of DMILO and DMILO-PGD in addressing common challenges in DM-based IP solvers.
@@ 1 Introduction
Compared to these conventional generative models like VAEs and GANs, diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019; Dhariwal & Nichol, 2021) have recently emerged as a powerful framework for modeling high-dimensional data distributions with remarkable fidelity. These models have demonstrated significant potential in addressing a variety of IPs, often achieving state-of-the-art (SOTA) performance. Motivated by the effectiveness of DMs in solving IPs, this work explores DM based IP solvers, aiming to further advance the field by leveraging the properties and strengths of DMs (Wang et al., 2023; Whang et al., 2022; Saharia et al., 2022b, a; Alkan et al., 2023; Cardoso et al., 2023; Chung et al., 2023a; Feng & Bouman, 2023; Rout et al., 2023; Wu et al., 2023; Aali et al., 2025; Chung et al., 2024; Dou & Song, 2024; Song et al., 2024; Wu et al., 2024; Sun et al., 2024).
@@ 1.1 Related Work
To address these issues, several methods employ projected gradient descent (PGD) (Shah & Hegde, 2018; Hyder et al., 2019; Peng et al., 2020; Liu & Han, 2022), using iterative projections to mitigate the risk of local minima. Some approaches allow for sparse deviations (Dhar et al., 2018) from the range of the generative model to capture signals outside the range. Additionally, the works (Daras et al., 2021, 2022) further perform optimization in intermediate layers to better align the reconstruction with the measurements.
@@
Based on how algorithms reconstruct the underlying signal $\bm{x}^{*}$ with prior distribution $p(\bm{x})$, works with pre-trained DMs can be further classified. Some methods apply Bayes’ rule (Fabian et al., 2023; Fei et al., 2023; Zhang et al., 2023) to decompose the conditional score $\nabla_{\bm{x}}\log p(\bm{x}|\bm{y})$ into a tractable unconditional score $\nabla_{\bm{x}}\log p(\bm{x})$ provided by the pretrained DMs and an intractable measurement-matching term, performing iterative corrections. For example, DDRM (Kawar et al., 2022) uses singular value decomposition (SVD) to approximate the measurement-matching term in spectral space. MCG (Chung et al., 2022) adapts Tweedie’s formula in linear IPs to approximate the reconstructed signal, followed by a gradient-based posterior correction. It also performs projection with a manifold constraint to ensure that corrections remain on the data manifold. DPS (Chung et al., 2023b) discards the projection step in the reverse process of MCG to prevent the samples from falling off the manifold in noisy tasks and generalizes to both linear and nonlinear tasks. $\Pi$GDM (Song et al., 2023) employs pseudo-inverse guidance to encourage data consistency between denoising results and the degraded image. Although these methods achieve high performance in many tasks, they are sometimes sensitive to guidance strength and may fail in simple scenarios, generating unnatural images or incorrect reconstructions.
@@ 2 Preliminaries
[…]
It is known from (Anderson et al., 1958) that the forward SDE in (2) has a corresponding reverse-time diffusion process evolving from $T$ down to 00. This reverse SDE can be written as (Ho et al., 2020; Song et al., 2021b):
@@
In practice, the score function $\nabla_{\bm{x}}\log p_{t}(\bm{x}_{t})$ is approximated by a neural network. A common approach is to train a noise prediction network $\bm{\epsilon}_{\theta}(\bm{x}_{t},t)$ to estimate the scaled score $-\sigma_{t}\,\nabla_{\bm{x}}\log p_{t}(\bm{x}_{t})$. The parameter set $\theta$ is optimized by minimizing the following objective:
@@
In the sampling process, DMs iteratively denoise an initial Gaussian noise vector to produce an image sample. Concretely, let $[\epsilon,T]$ be partitioned by $\epsilon=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, where $\epsilon>0$ is small to avoid numerical instabilities (Lu et al., 2022a). Then, by applying a first-order approximation to (8), the transition from $t_{i}$ to $t_{i-1}$ can be written as:
@@ 3.1 DMILO
1:  Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps $N$, noisy observation $\bm{y}$, forward model $\mathcal{A}(\cdot)$, iteration steps $J$, Lagrange multiplier $\lambda$
@@ 3.2 DMILO-PGD
1:  Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps $N$, noisy observation $\bm{y}$, forward model $\mathcal{A}(\cdot)$, iteration steps $E$, learning rate $\eta$, Lagrange multiplier $\lambda$
@@ 4 Theoretical Analysis
Let $\mathcal{S}\subseteq\mathbb{R}^{n}$. For some parameters $\gamma\in(0,1),\delta\geq 0$, a matrix $\bm{A}\in\mathbb{R}^{m\times n}$ is said to satisfy the S-REC$(\mathcal{S},\gamma,\delta)$ if for all $\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}$,
@@ 5.1 Inpainting and Super-resolution
For inpainting and super-resolution two linear tasks, we create evaluation sets by sampling 100 images each from CelebA (Liu et al., 2015), FFHQ (Karras et al., 2019), and LSUN-bedroom (Yu et al., 2015). [^4] The experimental results for LSUN-bedroom are presented in Appendix C due to the page limit. All images are resized to $256\times 256\times 3$ pixels. For super-resolution, we generate measurements by applying 4$\times$ bicubic downsampling, and for inpainting, we use a random mask with 70% missing pixels. Consistent with the setting in DMPlug (Wang et al., 2024), all measurements are corrupted by additive zero - mean Gaussian noise with a standard deviation of $\sigma=0.01$. We compare our methods with DDRM (Kawar et al., 2022), DPS (Chung et al., 2023b), $\Pi$GDM (Song et al., 2023), RED-diff (Song et al., 2023), DAPS (Zhang et al., 2025), DiffPIR (Zhu et al., 2023) and DMPlug (Wang et al., 2024).
@@
In the inpainting and super-resolution tasks, both DMILO and DMILO-PGD are configured with a Lagrange multiplier $\lambda=0.1$. For inpainting, during the optimization process, both DMILO and DMILO-PGD employ the Adam optimizer with an inner learning rate of 0.02. They run for 200 inner iterations and 5 outer iterations. In the case of DMILO-PGD, the outer learning rate $\eta$ is set to 0.5. For super-resolution, an inner learning rate of 0.02 is also used, along with 400 inner iterations and 10 outer iterations. For DMILO-PGD in super-resolution, the outer learning rate $\eta$ is set to 8. The quantitative results are presented in Table 2. Evidently, our DMILO-PGD approach surpasses all competing methods in most metrics. This clearly demonstrates its superior reconstruction performance, highlighting its effectiveness and competitiveness in these two linear IP tasks.
@@
Super-resolution (4×4$\times$4 ×) Inpainting (Random $70\%$)
@@
Ours vs. Best comp. $-0.047$ $+1.10$ $+0.018$ $+0.038$ $+0.06$ $-0.008$ $-0.028$ $+0.91$ $+0.017$ $-0.020$ $+0.46$ $+0.013$
@@ 5.2 Linear Image Deblurring
For linear image deblurring, we choose Gaussian deblurring and motion deblurring two linear image deblurring tasks. To validate the performance of different methods, we randomly sample 100 images from CelebA, FFHQ, and ImageNet (Deng et al., 2009). All images are resized to $256\times 256\times 3$ and the noise level $\sigma$ is set to $0.01$ similarly. We compare our methods with DPS, RED-diff, DPIR (Zhang et al., 2021) and DiffPIR, DMPlug. Note that we follow DMPlug to set the kernel size to $64\times 64$, which differs from the kernel size initially employed in DPIR and DiffPIR. Such a difference might impact the results. The results presented in Tables 3 and 4 indicates that our methods generally perform well in terms of both realism and reconstruction metrics, demonstrating its ability to effectively balance perceptual quality and distortion. They achieve top performance across all metrics for motion deblurring and achieve either the best or competitive results on Gaussian deblurring.
@@
Ours vs. Best comp. +4.36 -$0$0.41 -0.045 -14.06 -0.009 +0.54 0.008 -13.39 -0.008 -0.64 -0.035
@@
Ours vs. Best comp. -36.13 -0.073 +3.06 +0.082 -52.18 -0.093 +3.54 +0.089 -46.10 -0.184 +2.69 +0.083
@@ 5.3 Nonlinear Deblurring
For nonlinear deblurring, we use the learned blurring operators from (Tran et al., 2021) with a known Gaussian-shaped kernel and additive zero-mean Gaussian noise with a standard deviation of $\sigma=0.01$, following the setup of DMPlug. Similar to the process for linear tasks, our evaluation sets are created by sampling 100 images from CelebA, FFHQ, and LSUN. Each image is resized to $256\times 256\times 3$ pixels. As DDRM is only designed for linear IPs, we compare our DMILO and DMILO-PGD methods against DPS, $\Pi$GDM, RED-diff, and DMPlug.
@@
Ours vs. Best comp. $-0.037$ $+2.55$ $+0.038$ $-0.052$ $+2.65$ $+0.053$
@@ Acknowledgment
Then, based on Lemma A.1 and the well-established chaining arguments in (Bora et al., 2017; Liu & Scarlett, 2020a), we have that when $m=\Omega(k_{2}\log\frac{L_{1}n}{\delta}+k^{2}\log(3n))$, with probability $1-e^{-\Omega(m)}$, $\bm{A}$ satisfies S-REC$(\mathcal{X}-\mathcal{X},\gamma,\delta)$. Then, we have
@@
where (26) follows from $\bm{A}$ satisfies S-REC$(\mathcal{X}-\mathcal{X},\gamma,\delta)$, (28) follows from (18), and (29) follows from Lemma A.1 with setting $\epsilon=\frac{1}{4}$. This completes the proof.
@@ Appendix B Algorithms for BID
1:  Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps $N$, noisy observation $\bm{y}$, iteration steps $J$, Lagrange multiplier $\lambda$
@@
1:  Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps $N$, noisy observation $\bm{y}$, iteration steps $E$, learning rate $\eta_{x}$, $\eta_{k}$, Lagrange multiplier $\lambda$
@@ C.1 Results on LSUN-bedroom Dataset
Super-resolution (4×4$\times$4 ×) Inpainting (Random $70\%$) Nonlinear Deblurring
@@ C.4 Memory and Time Cost
DDRM DPS $\Pi$GDM RED-diff DiffPIR DAPS DMPlug DMPlug-Ckpt DMILO DMILO-PGD
@@ Appendix D Visualization of Experimental Results
Figure 7: Visualization of sample results for 4$\times$ super-resolution with additive Gaussian noise ($\sigma=0.01$).

### v3

@@ Abstract
Inverse problems (IPs) involve reconstructing signals from noisy observations. Recently, diffusion models (DMs) have emerged as a powerful framework for solving IPs, achieving remarkable reconstruction performance. However, existing DM-based methods frequently encounter issues such as heavy computational demands and suboptimal convergence. In this work, building upon the idea of the recent work DMPlug (Wang et al., 2024), we propose two novel methods, DMILO and DMILO-PGD, to address these challenges. Our first method, DMILO, employs intermediate layer optimization (ILO) to alleviate the memory burden inherent in DMPlug. Additionally, by introducing sparse deviations, we expand the range of DMs, enabling the exploration of underlying signals that may lie outside the range of the diffusion model. We further propose DMILO-PGD, which integrates ILO with projected gradient descent (PGD), thereby reducing the risk of suboptimal convergence. We provide an intuitive theoretical analysis of our approaches under appropriate conditions and validate their superiority through extensive experiments on diverse image datasets, encompassing both linear and nonlinear IPs. Our results demonstrate significant performance gains over state-of-the-art methods, highlighting the effectiveness of DMILO and DMILO-PGD in addressing common challenges in DM-based IP solvers. The code is available at https://github.com/StarNextDay/DMILO.git.
@@ 1 Introduction
Compared to these conventional generative models like VAEs and GANs, diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019; Dhariwal & Nichol, 2021) have recently emerged as a powerful framework for modeling high-dimensional data distributions with remarkable fidelity. These models have demonstrated significant potential in addressing a variety of IPs, often achieving state-of-the-art (SOTA) performance. Motivated by the effectiveness of DMs in solving IPs, this work explores DM based IP solvers, aiming to further advance the field by leveraging the properties and strengths of DMs (Wang et al., 2023; Whang et al., 2022; Saharia et al., 2022b, a; Alkan et al., 2023; Cardoso et al., 2023; Chung et al., 2023a; Feng & Bouman, 2023; Rout et al., 2023; Wu et al., 2023; Aali et al., 2025; Chung et al., 2024; Dou & Song, 2024; Song et al., 2024; Wu et al., 2024; Zhang et al., 2024; Sun et al., 2024; Janati et al., 2025; Zhang et al., 2025b).
@@ 1.1 Related Work
To address these issues, several methods employ projected gradient descent (PGD) (Shah & Hegde, 2018; Hyder et al., 2019; Peng et al., 2020; Liu & Han, 2022; Chen et al., 2025), using iterative projections to mitigate the risk of local minima. Some approaches allow for sparse deviations (Dhar et al., 2018) from the range of the generative model to capture signals outside the range. Additionally, the works (Daras et al., 2021, 2022) further perform optimization in intermediate layers to better align the reconstruction with the measurements.
@@
[…]
@@ 2 Preliminaries
$f(t)\;=\;\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\quad g^{2}(t)\;=\;\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}\;-\;2\,\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\,\sigma_{t}^{2}.$ (3)
It is known from (Anderson et al., 1958) that the forward SDE in (2) has a corresponding reverse-time diffusion process evolving from $T$ down to 0. This reverse SDE can be written as (Ho et al., 2020; Song et al., 2021b):
@@
In practice, the score function $\nabla_{\bm{x}}\log p_{t}(\bm{x}_{t})$ is approximated by a neural network. A common approach is to train a noise prediction network $\bm{\epsilon}_{\theta}(\bm{x}_{t},t)$ to estimate the scaled score $\sigma_{t}\,\nabla_{\bm{x}}\log p_{t}(\bm{x}_{t})$. The parameter set $\theta$ is optimized by minimizing the following objective:
@@
In the sampling process, DMs iteratively denoise an initial Gaussian noise vector to produce an image sample. Concretely, let $\epsilon,T]$ be partitioned by $\epsilon=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, where $\epsilon>0$ is small to avoid numerical instabilities (Lu et al., 2022a). Then, by applying a first-order approximation to (8), the transition from $t_{i}$ to $t_{i-1}$ can be written as:
@@ 3.1 DMILO
1:Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps NN, noisy observation $\bm{y}$, forward model $\mathcal{A}(\cdot)$, iteration steps $J$, Lagrange multiplier $\lambda$
@@ 3.2 DMILO-PGD
1:Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps NN, noisy observation $\bm{y}$, forward model $\mathcal{A}(\cdot)$, iteration steps $E$, learning rate $\eta$, Lagrange multiplier $\lambda$
@@ 4 Theoretical Analysis
[…]
@@ 5.1 Inpainting and Super-resolution
[…]
@@
In the inpainting and super-resolution tasks, both DMILO and DMILO-PGD are configured with a Lagrange multiplier $\lambda=0.1$. For inpainting, during the optimization process, both DMILO and DMILO-PGD employ the Adam optimizer with an inner learning rate of 0.02. They run for 200 inner iterations and 5 outer iterations. In the case of DMILO-PGD, the outer learning rate $\eta$ is set to 0.5. For super-resolution, an inner learning rate of 0.02 is also used, along with 400 inner iterations and 10 outer iterations. For DMILO-PGD in super-resolution, the outer learning rate $\eta$ is set to 8. The quantitative results are presented in Table 2. Our DMILO-PGD approach outperforms or achieves performance comparable to all competing methods across most evaluation metrics. This clearly demonstrates its superior reconstruction performance, highlighting its effectiveness and competitiveness in these two linear IP tasks.
@@
Super-resolution ($4\times$) Inpainting (Random $70\%$)
@@
Ours vs. Best comp. +0.004+0.004 +1.10+1.10 +0.018+0.018 +0.059+0.059 +0.03+0.03 0.008-0.008 0.000-0.000 +0.91+0.91 +0.012+0.012 +0.002+0.002 0.89-0.89 0.007-0.007
@@ 5.2 Linear Image Deblurring
For linear image deblurring, we choose Gaussian deblurring and motion deblurring two linear image deblurring tasks. To validate the performance of different methods, we randomly sample 100 images from CelebA, FFHQ, and ImageNet (Deng et al., 2009). All images are resized to $256\times 256\times 3$ and the noise level $\sigma$ is set to $0.01$ similarly. We compare our methods with DPS, RED-diff, DPIR (Zhang et al., 2021), DiffPIR, DMPlug, MGPS and DCPS. Note that we follow DMPlug to set the kernel size to $64\times 64$, which differs from the kernel size initially employed in DPIR and DiffPIR. Such a difference might impact the results. The results presented in Tables 3 and 4 indicates that our methods generally perform well in terms of both realism and reconstruction metrics, demonstrating its ability to effectively balance perceptual quality and distortion. They achieve top performance across most metrics for motion deblurring and achieve either the best or competitive results on Gaussian deblurring.
@@
Ours vs. Best comp. +12.54 +0.011 -0.41 -0.045 +26.19 +0.030 +0.54 -0.008 +79.46 +0.119 -0.64 -0.035
@@
Ours vs. Best comp. -4.11 -0.010 +2.97 +0.050 +3.93 -0.017 +3.54 +0.065 +10.29 -0.034 +2.00 +0.052
@@ 5.3 Nonlinear Deblurring
[…]
@@
Ours vs. Best comp. 0.037-0.037 +2.55+2.55 +0.038+0.038 0.052-0.052 +2.65+2.65 +0.053+0.053
@@ Acknowledgment
[…]
@@
[…]
@@ Appendix B Algorithms for BID
1:Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps NN, noisy observation $\bm{y}$, iteration steps $J$, Lagrange multiplier $\lambda$
@@
1:Input: Diffusion denoising model $\mathcal{G}(\cdot)=g_{1}\circ\cdots\circ g_{N-1}\circ g_{N}(\cdot)$, total number of sampling steps NN, noisy observation $\bm{y}$, iteration steps $E$, learning rate $\eta_{x}$, $\eta_{k}$, Lagrange multiplier $\lambda$
@@ C.1 Results on LSUN-bedroom Dataset
Super-resolution ($4\times$) Inpainting (Random $70\%$) Nonlinear Deblurring
@@ C.4 Memory and Time Cost
[…]
@@ Appendix D Visualization of Experimental Results
[…]